# TabulaX: Leveraging Large Language Models for Multi-Class Table Transformations

### v1

@@ Abstract
The integration of tabular data from diverse sources is often hindered by inconsistencies in formatting and representation, posing significant challenges for data analysts and personal digital assistants. Existing methods for automating tabular data transformations are limited in scope, often focusing on specific types of transformations or lacking interpretability. In this paper, we introduce TabulaX, a novel framework that leverages Large Language Models (LLMs) for multi-class tabular transformations. TabulaX first classifies input tables into four transformation classes—string-based, numerical, algorithmic, and general—and then applies tailored methods to generate human-interpretable transformation functions, such as numeric formulas or programming code. This approach enhances transparency and allows users to understand and modify the mappings. Through extensive experiments on real-world datasets from various domains, we demonstrate that TabulaX outperforms existing state-of-the-art approaches in terms of accuracy, supports a broader class of transformations, and generates interpretable transformations that can be efficiently applied.
Index Terms:
@@
Large Language Models, Heterogeneous Table Join, Data Integration, Tabular Data Transformation, Data Cleaning and Transformation.
I Introduction
@@
The rapid growth of publicly available data and increased reliance on third-party sources have underscored the need for efficient methods to combine and transform data from diverse sources. This is crucial for a wide range of users, including data analysts, scientists, and personal digital assistants. However, significant challenges arise due to variations in formatting and presentation across these sources. Even within a single organization, data such as names, addresses, and financial records may exist in multiple inconsistent formats. Target tables may be stored in formats such as spreadsheets, relational databases, web tables, and comma-separated files–common across organizations, the web, and public government datasets [1, 2].
@@
There has been a wide array of studies on automating the transformation of tabular data and ensuring consistency across diverse data sources [2, 3, 4, 5], yet significant challenges remain. Table I reports the performance, in terms of F1-score, of four state-of-the-art (SOTA) methods on KBWT [5], a dataset comprising tables extracted from a knowledge base (§V-A). Clearly, all SOTA models struggle and only GXJoin generates explainable transformations, while the others produce either an output string or a match decision, both of which lack interpretability. Figure 1 illustrates examples of source and target tables. In the first table pair, where individuals are identified by name in the source and username in the target, all SOTA models perform well. However, for the second pair, involving weight conversions between pounds and kilograms, and the third pair, linking company names with their CEOs, the models fail to produce meaningful results. While GPT-4o often produces correct outputs, reflected in its relatively high F1-Score in Table I, it does not yield transformations that are scalable to large datasets.
TABLE I: Performance of SOTA models on KBWT dataset
GPT-4o DTT [2] GXJoin [3] AFJ [6]
@@
Figure 1: Example input tables with formatting mismatch
@@
The problem can be formulated as given a small set of matched rows between source and target tables, we want to identify mappings that transform source values into target values, enabling an equi-join between the two tables. This paper explores two key research questions in this context: (1) Can higher performance be achieved in transforming tables that exhibit a wide range of syntactic and semantic patterns (as illustrated in Figure 1)? (2) Can the transformation process be made explainable? Aligned with prior work in the literature [2, 3, 7, 8], this work primarily focuses on transformations applied to key or join columns. However, the proposed approach has the potential to be extended to include additional columns or more complex settings.
@@
Challenges. There are a few critical challenges that must be addressed to tackle the problem. One major challenge is the vast multi-class search space due to the diverse nature of mismatches, as illustrated in Figure 1. These mismatches can range from string transformations, such as converting full names to usernames, to numeric conversions, like pounds to kilograms. In more complex cases, transformations may even require external knowledge bases (for example, linking a company name to its CEO). An effective solution must not only handle all of these classes but also navigate the vast search space of potential transformations within each class to identify the most appropriate mappings. Another challenge arises from the limited coverage of individual transformations. A single transformation may only be applicable to a subset of the rows, as demonstrated in the first table in Figure 1, where entries containing middle names are transformed differently than those without. A comprehensive framework must identify multiple transformations and develop rules to apply the appropriate transformation based on observed data patterns. Additionally, interpretability is a crucial requirement for many application areas, particularly those involving sensitive or high-stakes data. Users need to understand and verify the logic behind each transformation to ensure it meets their requirements. Furthermore, the ability to adjust or refine these mappings fosters trust and transparency in data integration processes. Finally, the issue of missing values in target tables presents an additional challenge. Ideally, we want to impute missing values using corresponding source data, ensuring that the integration process remains comprehensive and reliable, even in the presence of incomplete information.
@@
Limitations of existing approaches. Existing methods for addressing table transformations can be broadly categorized into (1) methods that learn a matching function for a given pair of values or records, and (2) methods that generate mapping rules between source and target tables [2, 3, 9, 7, 8]. The first group includes techniques focused on learning similarity functions [6, 10] or mapping functions [11, 12], as well as approaches for matching entity records that refer to the same real-world objects but differ in format or representation [13, 14, 15]. These approaches typically rely on machine learning models or (semi-)manually defined similarity metrics to rank and select the best candidate for each corresponding value. While effective for addressing formatting mismatches in joining tabular data, they lack the ability to generate explicit mappings to transform the source formatting into the target. This limits their flexibility in modifying transformations, may lead to non-interpretable functions, and restricts their use primarily to fuzzy joins. The second group of approaches focuses on generating mapping rules. Many of these studies are limited to string-based transformations and rely on an exhaustive search of the parameter space to find the appropriate mapping. Some methods are constrained to using a single transformation function per table [7, 16, 8, 17], while others depend entirely on retrieving functions from external sources [4, 18]. Also, some approaches employ pretrained language models, which return mappings that are not interpretable [2].
@@
Our approach. We introduce TabulaX, a novel data integration framework that leverages Large Language Models (LLMs) for multi-class tabular transformations. Similar to existing approaches, TabulaX transforms the entity or join column in the source table to its corresponding representation in the target. However, the mapping functions generated by our framework cover a much broader range of transformations, including textual, numeric, algorithmic, and those requiring external knowledge. A key feature of TabulaX lies in the interpretability and usability of its transformations, which are generated as numeric functions or programming language code, unlike some approaches in the literature that rely on their own domain specific languages [16, 8, 17]. Our experimental evaluation reveals that TabulaX outperforms existing state-of-the-art approaches in terms of accuracy, supports a broader class of transformations, and delivers mappings that can be efficiently applied to large-scale tables and datasets.
@@
Contributions. Our contributions can be summarized as follows: (1) We propose TabulaX, a new framework for Multi-class Tabular Transformations, capable of handling a broad range of transformations—including textual, numeric, algorithmic, and those requiring external knowledge. (2) We introduce a classification mechanism that determines the transformation class of input tables, enabling TabulaX to apply appropriate methods for diverse data mismatches effectively. (3) We develop methods to generate human-interpretable transformation functions such as numeric formulas and programming language code, enhancing transparency and allowing users to understand and modify the mappings. (4) Through extensive experiments on real-world datasets from various domains, we demonstrate that TabulaX outperforms existing state-of-the-art approaches in terms of accuracy and supports a broader class of transformations. (5) We make all our resources publicly available, including our code, framework, and datasets, to facilitate further research [^1] https://github.com/arashdn/TabulaX.
II Related Works
Related work can be categorized into example-driven tabular transformations for join and leveraging LLMs for tabular data processing.
II-A Example-Driven Tabular Data Transformation for Join
This category closely aligns with our work and has been an active area of research in recent years [2, 9, 7, 16, 8, 17]. Early approaches such as FlashFill [8] and BlinkFill [16], particularly effective for spreadsheet data, generate substring-based transformations by constructing an input graph from user-provided examples and traversing the graph to map source values to their target representations. However, these approaches heavily rely on the accuracy of the examples and struggle when dealing with noisy data. Zhu et al. introduced Auto-join [7], which employs predefined string-based transformation units, such as substring and split, to generate mappings. Auto-join automatically generates examples using token matching and handles noise by partitioning input examples into subsets, applying a recursive backtracking algorithm to find the best transformation for each subset, and ranking them accordingly. While this approach improves noise handling and reduces the search space by using pre-defined transformation units, its backtracking algorithm can be computationally intensive and may struggle with high levels of noise.
To address the aforementioned limitations, Nobari et al. developed the Common String-based Transformer (CST) [9], which further reduces the search space for string-based transformations by leveraging common text sequences between source and target examples as textual evidence to form the transformation skeleton. CST uses the same string-based transformation units as Auto-join but generates transformations independently for each row to better handle input noise. These row-level transformations are then ranked based on their coverage to form the final transformation set. Building on CST, GXJoin [3] redefined the transformation units to provide a more general set of mappings, improving both coverage and robustness to noise.
CST and GXJoin offer improved noise handling and runtime efficiency compared to earlier methods but remain limited to a narrow set of predefined substring-based transformations. These methods are effective only when there are long matching sequences between the source and target examples. To address these constraints, a more recent study, referred to as DTT, leverages a fine-tuned Language Model (LM) to transform source values into target representations. This approach eliminates the need for exhaustive parameter-space searches, which often relies on pruning techniques that can miss some transformations. Unlike previous methods, DTT is not limited to predefined functions and is trained on a large set of textual transformations. While DTT overcomes many limitations of function-based transformers, it is still trained exclusively on a synthetic dataset of string-based transformations, and its mapping space is a non-interpretable latent space. Our work aims to tackle these problems by introducing a novel class-aware framework exploiting LLMs to generate interpretable mappings.
II-B Large Language Models for processing Tabular Data
The widespread adoption of first-generation pretrained language models such as T5 [19], BERT [20], GPT-2 [21], and GPT-3 [22] has led to a variety of such models being employed to table embedding and understanding [23, 24, 25]. Models such as TaBERT [26], TURL [27], and TABBIE [28] use an encoder-only structure, making them better suited for discriminative tasks like entity matching [29, 13] and question answering [30, 26, 28]. Other models, like RPT [31], use an encoder-decoder architecture, which makes them more suitable for generative tasks such as data to text [32, 33, 34]. This group of LMs typically have less than 500 million parameters and rely on fine-tuning techniques to adapt to various tasks. The aforementioned models are trained on tabular data and may require further fine-tuning for specific tasks. For instance, DTT [2] is a framework based on a fine-tuned version of ByT5 [35], developed to transform tables for joinability.
More recently, the large language models with billions of parameters such as ChatGPT [^2] https://chatgpt.com, Gemini [^3] https://gemini.google.com, and Llama 3 [36] have outperformed earlier generations of language models across various tasks. LLMs are typically general-purpose and can be adapted to many tasks by providing only a few high-quality examples (i.e., shots) in the prompt, without the need for further fine-tuning [22]. This has inspired a growing body of research on prompt-tuning techniques [37], such as Self-consistency [38], which generates diverse reasoning chains by sampling from the model’s output instead of using a greedy approach, and Chain-of-Thought Prompting [39], where problems are broken down into smaller, manageable steps, encouraging the model to articulate its thought process step-by-step. This method has been shown to enhance the model’s ability to tackle complex problems, rather than simply asking for a direct answer.
A number of studies have explored the use of LLMs for tabular data through supervised fine-tuning [40, 41] and prompt-tuning approaches [42]. When leveraging language models to process tabular data, it is crucial to ensure that the table is properly serialized and prompted [2, 43]. Several methods have been proposed to transform tables into sequences of tokens [2, 42, 31, 28], while preserving the structural relationships necessary for downstream tasks. These approaches are orthogonal to ours, with pre-processing the tables to prepare them to be prompted to a LLM and exploiting LLMs to process tabular data.
III Multi-Class Transformation Discovery
@@
Our aim is to transform the values in a source table column into their corresponding values in a target column, guided by a small set of provided examples. We assume that both the source values to be transformed and user-provided examples are available, which helps narrow the scope to concentrate specifically on data transformation tasks [2, 9]. When such examples are not provided by users, methods such as unequal joining [6, 11, 12] or token-based example generation [9, 7] can be employed to produce a set of examples. However, these automatically generated examples may include noise or invalid matches.
@@
$f:S\rightarrow T|\forall s_{j}\in S^{\prime}((s_{j},f(s_{j})\in E.$ (1)
@@
Figure 2: The architecture of our framework
@@
The majority of existing approaches in the literature limit their scope when addressing this problem. Some methods focus only on string-based transformations [2, 7, 8], others only handle entity matching [15, 29], and some retrieve mapping functions from large code repositories [4, 18]. One may argue that recent commercially available LLMs, such as GPT and Gemini, could address these challenges. To evaluate the performance of LLMs in this context, we conducted an experiment using the GPT model [^4] We used the gpt-4o-2024-05-13 model from the OpenAI API for this experiment on our benchmark consisting of diverse real-world tables. However, the out-of-the-box model struggled to recognize many textual patterns, and frequently produced hallucinated or irrelevant outputs in cases requiring external knowledge. Moreover, LLMs are not expected to handle numerical data well [44], and our experiments also confirmed that the model was unable to perform numerical transformations effectively.
@@
  • Numerical: This class encompasses input tables where both source and target values are numerical, such as in the conversion from pounds to kilograms illustrated in Figure 1. It is important to note that nominal numbers, such as postal codes or ISBNs, do not fall in this category.
  • Algorithmic: This class includes input tables where transformations can be carried out using algorithms that do not rely on external knowledge sources and do not fall under the string-based or numerical classes. Examples include transformation such as Gregorian to Hijri dates, characters to ASCII codes, or binary to hexadecimal numbers.

### v2

@@ Abstract.
The integration of tabular data from diverse sources is often hindered by inconsistencies in formatting and representation, posing significant challenges for data analysts and personal digital assistants. Existing methods for automating tabular data transformations are limited in scope, often focusing on specific types of transformations or lacking interpretability. In this paper, we introduce TabulaX, a novel framework that leverages Large Language Models (LLMs) for multi-class column-level tabular transformations. TabulaX first classifies input columns into four transformation types—string-based, numerical, algorithmic, and general—and then applies tailored methods to generate human-interpretable transformation functions, such as numeric formulas or programming code. This approach enhances transparency and allows users to understand and modify the mappings. Through extensive experiments on real-world datasets from various domains, we demonstrate that TabulaX outperforms existing state-of-the-art approaches in terms of accuracy, supports a broader class of transformations, and generates interpretable transformations that can be efficiently applied.
@@
Large Language Models, Heterogeneous Table Join, Data Integration, Data Transformation, Data Cleaning and Transformation
PVLDB Reference Format:
PVLDB, 18(11): , 2025.
This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 11 ISSN 2150-8097.
doi:ArXiv Version
1. Introduction
@@
The rapid growth of publicly available data and increased reliance on third-party sources have underscored the need for efficient methods to combine and transform data from diverse sources. This is crucial for a wide range of users, including data analysts, scientists, and personal digital assistants. However, significant challenges arise due to variations in formatting and presentation across these sources. Even within a single organization, data such as names, addresses, and financial records may exist in multiple inconsistent formats. Target tables may be stored in formats such as spreadsheets, relational databases, web tables, and comma-separated files—common across organizations, the web, and public government datasets (Miller et al., 2018; Dargahi Nobari and Rafiei, 2024).
@@
There has been a wide array of studies on automating the transformation of tabular data and ensuring consistency across diverse data sources (Dargahi Nobari and Rafiei, 2024; Omidvartehrani et al., 2024; He et al., 2018a; Abedjan et al., 2016), yet significant challenges remain. Table 1 reports the performance, in terms of F1-score, of four state-of-the-art (SOTA) methods on KBWT (Abedjan et al., 2016), a dataset comprising tables extracted from a knowledge base (§5.1). Clearly, all SOTA models struggle and only GXJoin generates explainable transformations, while the others produce either an output string or a match decision, both of which lack interpretability. Figure 1 illustrates examples of source and target tables. In the first table pair, where individuals are identified by name in the source and username in the target, all SOTA models perform well. However, for the second pair, involving weight conversions between pounds and kilograms, and the third pair, linking company names with their CEOs, the models fail to produce meaningful results. While GPT-4o often produces correct outputs, reflected in its relatively high F1-Score in Table 1, it does not yield transformations that are scalable to large datasets.
Table 1. Performance of SOTA models on KBWT dataset
GPT-4o DTT (Dargahi Nobari and Rafiei, 2024) GXJoin (Omidvartehrani et al., 2024) AFJ (Li et al., 2021)
@@
Figure 1. Example input tables with formatting mismatch
@@
The problem can be formulated as given a small set of matched rows between source and target tables, we want to identify mappings that transform source values into target values, enabling an equi-join between the two tables. This paper explores two key research questions in this context: (1) Can higher performance be achieved in transforming tables that exhibit a wide range of syntactic and semantic patterns (as illustrated in Figure 1)? (2) Can the transformation process be made explainable? Aligned with prior work in the literature (Dargahi Nobari and Rafiei, 2024; Omidvartehrani et al., 2024; Zhu et al., 2017; Gulwani, 2011), this work primarily focuses on transformations applied to key or join columns. However, the proposed approach has the potential to be extended to include additional columns or more complex settings.
@@
Challenges. There are a few critical challenges that must be addressed to tackle the problem. One major challenge is the vast multi-class search space due to the diverse nature of mismatches, as illustrated in Figure 1. These mismatches can range from string transformations, such as converting full names to usernames, to numeric conversions, like pounds to kilograms. In more complex cases, transformations may even require external knowledge bases (for example, linking a company name to its CEO). An effective solution must not only handle all of these classes but also navigate the vast search space of potential transformations within each class to identify the most appropriate mappings. Another challenge arises from the limited coverage of individual transformations. A single transformation may only be applicable to a subset of the rows, as demonstrated in the first table in Figure 1, where entries containing middle names are transformed differently than those without. A comprehensive framework must identify multiple transformations and develop rules to apply the appropriate transformation based on observed data patterns. Additionally, interpretability is a crucial requirement for many application areas, particularly those involving sensitive or high-stakes data. Users need to understand and verify the logic behind each transformation to ensure it meets their requirements. Furthermore, the ability to adjust or refine these mappings fosters trust and transparency in data integration processes.
@@
Limitations of existing approaches. Existing methods for addressing table transformations can be broadly categorized into (1) methods that learn a matching function for a given pair of values or records, and (2) methods that generate mapping rules between source and target tables (Dargahi Nobari and Rafiei, 2024; Omidvartehrani et al., 2024; Dargahi Nobari and Rafiei, 2022; Zhu et al., 2017; Gulwani, 2011). The first group includes techniques focused on learning similarity functions (Li et al., 2021; Deng et al., 2014) or mapping functions (He et al., 2015; Wang and He, 2017), as well as approaches for matching entity records that refer to the same real-world objects but differ in format or representation (Akbarian Rastaghi et al., 2022; Li et al., 2020a; Zhao and He, 2019). These approaches typically rely on machine learning models or (semi-)manually defined similarity metrics to rank and select the best candidate for each corresponding value. While effective for addressing formatting mismatches in joining tabular data, they lack the ability to generate explicit mappings to transform the source formatting into the target. This limits their flexibility in modifying transformations, may lead to non-interpretable functions, and restricts their use primarily to fuzzy joins. The second group of approaches focuses on generating mapping rules. Many of these studies are limited to string-based transformations and rely on an exhaustive search of the parameter space to find the appropriate mapping. Some methods are constrained to using a single transformation function per table (Zhu et al., 2017; Singh, 2016; Gulwani, 2011; Gulwani et al., 2012), while others depend entirely on retrieving functions from external sources (He et al., 2018a, b). Also, some approaches employ pretrained language models, which return mappings that are not interpretable (Dargahi Nobari and Rafiei, 2024).
@@
Our approach. We introduce TabulaX, a novel data integration framework that leverages Large Language Models (LLMs) for multi-class column-level tabular transformations. Similar to existing approaches, TabulaX transforms the entity or join column in the source table to its corresponding representation in the target. However, the mapping functions generated by our framework cover a much broader range of transformations, including textual, numeric, algorithmic, and those requiring external knowledge. A key feature of TabulaX lies in the interpretability and usability of its transformations, which are generated as numeric functions or programming language code, unlike some approaches in the literature that rely on their own domain specific languages (Singh, 2016; Gulwani, 2011; Gulwani et al., 2012). Our experimental evaluation reveals that TabulaX outperforms existing state-of-the-art approaches in terms of accuracy, supports a broader class of transformations, and delivers mappings that can be efficiently applied to large-scale tables and datasets.
@@
Contributions. Our contributions can be summarized as follows: (1) We propose TabulaX, a new framework for Multi-class Tabular Transformations, capable of handling a broad range of transformations—including textual, numeric, algorithmic, and those requiring external knowledge. (2) We introduce a classification mechanism that enables TabulaX to effectively apply tailored transformation strategies to diverse types of data mismatches. (3) We develop methods for generating human-interpretable transformation functions, such as numeric formulas and executable code, which improve transparency and allow users to understand, verify, and modify the mappings. (4) Through extensive experiments on real-world datasets from various domains, we demonstrate that TabulaX outperforms existing state-of-the-art approaches in terms of accuracy and supports a broader class of transformations. (5) We publicly release the TabulaX framework, source code, and datasets to support reproducability and foster further research [^1] https://github.com/arashdn/TabulaX.
2. Related Works
We organize related work into three main areas: (1) example-driven table transformations, (2) the use of large language models for tabular data, and (3) techniques for integrating tabular datasets.
2.1. Example-Driven Table Transformations
This line of work, closely aligns with ours and has been an active area of research in recent years (Dargahi Nobari and Rafiei, 2024, 2022; Zhu et al., 2017; Singh, 2016; Gulwani, 2011; Gulwani et al., 2012). Early approaches such as FlashFill (Gulwani, 2011) and BlinkFill (Singh, 2016) targeted spreadsheet data using substring-based transformations derived from user examples. Auto-join (Zhu et al., 2017) improves robustness by employing predefined string transformation units and a recursive backtracking algorithm to partition and transform noisy input. CST (Dargahi Nobari and Rafiei, 2022) builds on this by extracting common text fragments as transformation skeletons, applying row-level transformations, and ranking them by coverage. GXJoin (Omidvartehrani et al., 2024) generalizes these transformation units, improving both coverage and noise tolerance.
While CST and GXJoin improve efficiency and noise handling, they remain limited to a narrow set of predefined substring-based operations and require long textual overlaps. To address these constraints, DTT (Dargahi Nobari and Rafiei, 2024) introduces a fine-tuned language model for example-driven transformations. This approach eliminates the need for predefined functions or pruning-based searches but relies on synthetic training data and produces transformations in a non-interpretable latent space. Our work addresses these problems by introducing a novel class-aware LLM-based framework capable of generating interpretable, human-readable transformation functions, bridging expressiveness, generalization, and transparency.
Some structural transformation methods, such as Foofah (Jin et al., 2017) and Explain-Da-V (Shraga and Miller, 2023), focus on reshaping entire tables. While they target cell- and grid-level alignment for schema matching, our column-level mappings can complement these approaches by handling diverse value-level mismatches in integration tasks.
2.2. Large Language Models for Tabular Data
Early pretrained models such as T5 (Raffel et al., 2020), BERT (Devlin et al., 2019), and GPT-3 (Brown et al., 2020) have inspired the development of tabular-specific LLMs such as TaBERT (Yin et al., 2020), TURL (Deng et al., 2020), and TABBIE (Iida et al., 2021). These encoder-based models are better suited for discriminative tasks like entity matching (Li et al., 2020b; Akbarian Rastaghi et al., 2022) and QA (Chen et al., 2020; Yin et al., 2020; Iida et al., 2021), while models like RPT (Tang et al., 2021) and ByT5 (Xue et al., 2022) support generative tasks such as data to text (Parikh et al., 2020; Kale and Rastogi, 2020; Thorne et al., 2021). Tabular LMs typically have less than 500M parameters and rely on fine-tuning for task adaptation. DTT (Dargahi Nobari and Rafiei, 2024), for instance, is built on fine-tuned ByT5 (Xue et al., 2022) for table joinability.
Recent advances in LLMs such as ChatGPT [^2] https://chatgpt.com, Gemini [^3] https://gemini.google.com, and Llama 3 (Dubey et al., 2024) have enabled prompt-based methods that require minimal or no fine-tuning. Techniques such as self-consistency (Wang et al., 2022) and chain-of-thought prompting (Wei et al., 2022) improve LLM reasoning and performance by guiding generation through structured prompts. Some studies apply LLMs to tabular tasks using supervised fine-tuning (Zhang et al., 2023; Li et al., 2024) and prompt-based approaches (Sui et al., 2024), emphasizing the importance of table serialization (Dargahi Nobari and Rafiei, 2024; Nahid and Rafiei, 2024; Sui et al., 2024). These approaches are complementary to ours, improving how LLMs process and reason about tabular data.
2.3. Tabular Data Integration
Research focused on linking and integrating structured datasets is closely related to our goals. A comprehensive overview of foundational methods is provided by Doan et al. (Doan et al., 2012), while more recent developments, particularly in the context of data lakes and open data integration, are reviewed by Miller et al. (Miller, 2018) and Khatiwada et al. (Khatiwada et al., 2022). Our work also connects with efforts to detect linkage points across datasets (Hassanzadeh et al., 2013). Notably, when schema elements (e.g., dbpedia:stockSymbol and dbpedia:stockTicker) are interpreted as cell values, their alignment can be viewed as a form of cell-level transformation. This perspective supports the idea that schema-level and value-level integration can be unified under a single transformation framework, which our method aims to achieve through class-aware, interpretable mappings.
3. Multi-Class Transformation Discovery
@@
Our aim is to transform the values in a source table column into their corresponding values in a target column, guided by a small set of provided examples. We assume that both the source values to be transformed and user-provided examples are available, which helps narrow the scope to concentrate specifically on data transformation tasks (Dargahi Nobari and Rafiei, 2024, 2022). When such examples are not provided by users, methods such as unequal joining (Li et al., 2021; He et al., 2015; Wang and He, 2017) or token-based example generation (Dargahi Nobari and Rafiei, 2022; Zhu et al., 2017) can be employed to produce a set of examples, with the caveat that these automatically generated examples may include noise or invalid matches.
@@
(1) $f:S\rightarrow T|\forall s_{j}\in S^{\prime}((s_{j},f(s_{j})\in E.$
@@
Figure 2. The architecture of our framework
@@
Most existing approaches in the literature limit their scope when addressing this problem. Some methods focus only on string-based transformations (Dargahi Nobari and Rafiei, 2024; Zhu et al., 2017; Gulwani, 2011), others only handle entity matching (Zhao and He, 2019; Li et al., 2020b), and some retrieve mapping functions from large code repositories (He et al., 2018a, b). One may argue that recent commercially available LLMs, such as GPT and Gemini, could address these challenges. To evaluate the performance of LLMs in this context, we conducted an experiment using the GPT model [^4] We used the gpt-4o-2024-05-13 model from the OpenAI API for this experiment on our benchmark consisting of diverse real-world tables. However, the out-of-the-box model struggled to recognize many textual patterns, and frequently produced hallucinated or irrelevant outputs in cases requiring external knowledge. Moreover, LLMs are not expected to handle numerical data well (Chang et al., 2024), and our experiments also confirmed that the model was unable to perform numerical transformations effectively.
@@
  • Numerical: This class encompasses input tables where both source and target values are rational numbers and exhibit an underlying mathematical relationship. An example is the conversion from pounds to kilograms illustrated in Figure 1. It is important to note that nominal numbers, such as postal codes or ISBNs, do not fall in this category.
  • Algorithmic: This class includes input tables where an algorithm exists to transform source values into target values, but the transformation goes beyond numerical functions or sequences of string operations. Examples include transformation such as Gregorian to Hijri dates, characters to ASCII codes, or binary to hexadecimal numbers.